# NeWRF: A Deep Learning Framework for Wireless Radiation Field Reconstruction and Channel Prediction

### v1

@@ 4.3 Improving the Convergence of NeWRF
One key difference between wireless measurements and images is that wireless antennas capture signals from all directions. Thus, in wireless, finding the ray direction to trace is a non-trivial problem. One potential solution is to search exhaustively along a discrete set of azimuth and elevation angle combinations that span all directions. However, this approach has two fatal drawbacks. First, the actual ray directions lie in a continuous space. Approximating them with a discrete set of ray directions would require an extremely dense grid of angles, which significantly increases the memory consumption for training the model. Secondly, for each receiver location, there are only 10$\sim$20 ray directions contributing significantly to the channel measurement, whereas a search grid of ${}^{\circ}$ resolution involves $360\times 180=64,800$ ray angles. As a result, optimizing the model to find the actual ray directions from such a large ray space is very difficult, if not impossible. To solve this issue, we leverage the fact that the Direction-of-Arrivals (DoA) of wireless signals can be measured using an antenna array at the receiver location Xiong and Jamieson (2013)  [^2] For readers interested in how antenna arrays work, we refer to Appendix C for a detailed description.. We use DoA measurements as prior knowledge and let the learning algorithm trace rays primarily toward directions identified by the DoA measurements. Through our experiments, we find that this DoA-guided ray-casting approach significantly reduces the training time and memory consumption, as well as enables the learning algorithm to converge quickly.
@@ 4.4 Optimizing NeWRF
We simultaneously optimize two MLP models for each environment. One "coarse" and one "fine". We use mini-batch gradient descent to train our models where at each iteration, we randomly pick a batch of receiver locations from the training set and obtain the ground truth channel and DoA measurements. We cast rays toward each DoA and use the stratified sampling strategy Mildenhall et al. (2021) to take coarse-grained samples uniformly distributed on each ray. We use these samples to query the "coarse" model and use the procedures described in Section 4.2 to synthesize coarse predictions of the channel for each receiver location. We use the predicted weights $w_{\mathbf{p}}$ of coarse samples to perform a round of hierarchical sampling which densely samples regions with high weight values. We query the "fine" model with the union of coarse and fine samples and synthesize the fine channel predictions. The coarse and fine channel predictions are each compared with the ground truth to calculate loss and update the parameters of each model.
@@
$\vspace{-10pt}L(\mathbf{h},\hat{\mathbf{h}})=\frac{\sum|\hat{\mathbf{h}}-\mathbf{h}|^{2}}{\sum|\mathbf{h}|^{2}}$ (8)
@@
Further details about our training pipeline and model architecture can be found in Appendix A.
@@ 4.5 NeWRF’s Scene Representation
Figure 6: The nature of wireless scenes: While backtracing direct rays leads to the transmitter, backtracing reflected rays consistently intersects at the virtual images of the transmitter.
We train the model as described in the previous section and find it has good generalization performance. However, we find that the model learns a representation of the wireless radiation scene that was not originally expected. Despite that, the scene still allows the model to predict wireless fields (channels) even in room-scale environments with merely a sparse set of channel measurements. In this section, we present our findings and more insights into it.
@@
To better understand why this is the case, we illustrate a simple wireless scene with a single transmitter and two reflective surfaces in Figure 6. Three receivers are placed in random locations to measure the wireless channel. The signals emitted by the transmitter propagate via three major paths to reach each receiver: one direct line-of-sight path and two reflection paths  [^3] We omit the second order reflection here for simplicity; however it can also be characterized in the same manner.. Since the receivers have no notion about the environment, they simply trace the rays back in the direction from which they arrived. Because of the law of reflection, the rays traced by all the receivers intersect at the locations of the virtual images of the transmitter and the transmitter itself. As a result, these points are sampled more frequently than other points in the space, giving the model an inductive bias to put higher volume densities at these locations. This ultimately allows the model to learn to characterize the effect of a reflector on the amplitude and phase of its incoming signal.
We note that this scene is one of many possible solutions to this optimization problem of assigning values to the samples along the rays such that the synthesized channel has a minimum difference from the ground truth channels in the training set. However, the model seems to have such a bias that it is more likely to converge to this representation of the scene. Although this representation is not what we had initially expected, it does reveal the simple nature of radiation scenes and allow for accurate channel synthesis from sparse measurements.
@@ 4.6 Predicting Channels at Unvisited Locations
While DoA information is known for our training data, it is not for the test data. Because of this, a new ray-casting technique was required for inference time predictions. Simply casting rays across a grid of all directions is problematic. Depending on the resolution of angles used for sampling we could under or overestimate the contribution of a virtual transmitter on a channel. When the resolution is too low, virtual transmitters could be missed and when it is too high, virtual transmitters could be sampled multiple times and contribute too much to the channel. For this reason, we developed a novel unsupervised learning-based ray-searching algorithm that leverages our discovery of the nature of wireless scenes to extract DoA information and thereby intelligently cast rays at new locations.
Our ray-searching algorithm consists of six steps:
@@
    Initialization: Initialize a set of rays $\mathbb{Q}$ to be densely pointing in all directions.
@@
    Querying: Sample all of the rays in $\mathbb{Q}$ and query the MLP for volume density predictions.
@@
    Pruning: Calculate the cumulative volume densities by summing the volume density predictions along each ray. Remove the rays whose cumulative density are lower than a threshold $t$ from the set $\mathbb{Q}$. We use $t=30$ in practice.
@@
    Merging: Within the remaining rays in $\mathbb{Q}$, to avoid oversampling the same (virtual) transmitter, we use the DBSCAN Ester et al. (1996) algorithm to cluster the rays based on directions and merge the rays within the same cluster to obtain a merged ray set $\mathbb{M}$.
@@
    Expanding: After pruning and merging, the rays in set $\mathbb{M}$ are very close to the ground truth ray, however, they may still be off by a small margin which can reduce accuracy. Therefore, we further refine the ray directions by emitting more rays near each ray in $\mathbb{M}$ with a small delta. The expanded ray set is named $\mathbb{X}$.
  6. 6.
    Iterating: Assign $\mathbb{Q}:=\mathbb{X}$ and Repeat step 2 to 5 until convergence.
We note that this algorithm can potentially discover propagation paths that are blocked. To solve this issue, we further filter the discovered paths using the DoA measurements of the nearest receiver location in the training set.
@@ 5 Evaluation
Datasets: We generate simulation datasets for evaluation using Wireless Insite Remcom (2023), a commercial wireless ray tracing simulation software, and MATLAB. Both Wireless Insite and MATLAB can run ray-tracing simulations. Wireless Insite also includes effects such as diffraction and penetration to support more realistic channel simulation. We use both simulators for cross-validation. We create datasets for three 3D environment models MATLAB (2023a, b); Free3D (2023). Figure 7 shows the geometry and dimensions of the three models. We put a transmitter in each environment and randomly spread multiple receivers throughout the space. We collect 443, 975, and 1907 measurements in each environment, respectively. We use 80% of the measurements for training the model and the remaining 20% for testing.
Figure 8: Channel Prediction Results: a) The predicted channel for the test set. Each dot represents the channel at one test location. b) The predicted channel components for a single location. Each dot represents a single propagation path. The arrow exhibits the sum of all paths; c) The channel prediction accuracy in different environments compared with baseline methods. d), e), and f) channel prediction accuracy for different training set sample densities, environment materials, and signal frequencies. Perfect Reflector is an ideal material that reflects all signals without any attenuation. Signal frequencies are selected as common frequency bands for various wireless technologies, such as WiFi, Bluetooth, RFID.
@@ 5.1 Channel Prediction Results
Figure 7(a) provides a visualization of the predicted channel for the test set measurements. Since the wireless channels are represented as complex numbers, we plot the real and imaginary parts of the channels in the x and y-axis, respectively. Each dot in the plot represents the channel for a particular receiver location. We plot the ground truth in red, and NeWRF’s prediction in blue. Our predictions align very well with the ground truth. Moreover, NeWRF can even predict individual path components for a particular receiver location as shown in Figure 7(b). In this plot, each dot represents one propagation path of the wireless signal; the arrow indicates the sum of all the components. The predicted channel components align very well with the ground truth, which indicates that NeWRF can predict the channel components accurately.
@@
3) NeRF2: We use the open-sourced code of NeRF2 Zhao (2023) and train it on our datasets. We swap the role of transmitters and receivers in our dataset to accommodate their settings.
We use Signal-to-Noise Ratio (SNR), as our evaluation metric, which is defined as the SNR = $-10\log_{10}(\text{NMSE})$, where NMSE is Normalised Mean Square Error and defined in Eq (8). Higher SNR indicates a higher prediction accuracy. For example, a 20 dB SNR is equivalent to a 1% relative squared error in the prediction. Figure 7(c) shows the results for three environments. NeWRF outperforms all of the baseline methods in all environments. Note, NeRF2 does not perform well on the datasets since the measurement density (0.4 $measurements/ft^{3}$) of our datasets is much lower than what they require (178 $measurement/ft^{3}$). Finally, our results show that although the performance of NeWRF degrades as the environment becomes larger and more complex (such as a large office space with lots of furniture), it still achieves more than 5 dB SNR which is sufficient for wireless applications.
@@ 5.2 Measurement Density of Training Set
We show the effect of training set sample density on the performance of NeWRF. To demonstrate this, we use the conference room dataset. In this experiment, we use 25%, 50%, 75%, and 100% of the training set measurements to train our model, which corresponds to 0.1, 0.2, 0.3, and 0.4 $measurements/ft^{3}$, respectively. We evaluate the trained model on the same test set for a fair comparison. As is shown in Figure 7(d), NeWRF can effectively learn the wireless radiation scene, and predict the channel accurately with sample density as low as 0.2 $measurements/ft^{3}$, which is around 800$\times$ lower than that of NeRF2Zhao et al. (2023). This would significantly reduce the cost of wireless site surveys and improve the capability of dead spot detection.
@@ 5.3 Impact of Materials Types and Signal Frequencies
The attenuation and phase rotation of wireless signals during reflection are largely determined by the material that constitutes the reflector. In particular, the material’s electrical properties, such as permittivity and conductivity, play significant roles. As such, the wireless properties of an environment can change greatly as the materials it is made of change. Furthermore, even the same materials have different behaviors when they interact with signals of different frequencies. NeWRF is able to characterize both of these effects, as shown in Figure 7(e) and 7(f), where we train a separate model for each material type and signal frequency. NeWRF maintains high prediction accuracy in all these scenarios.
@@ Appendix A Implementation Details
We implement NeWRF in Pytorch. Figure 10 shows the training pipeline. We use a batch size of 32 for each iteration of training. As in NeRF, we use positional encoding to map the input coordinates to higher dimensional space to facilitate the model to learn high-frequency variation of the wireless scene Mildenhall et al. [2021]. We use 10 frequencies to encode the spatial coordinates $(x,y,z)$, and 4 frequencies for the direction coordinates.
Figure 10 shows the MLP model architecture of NeWRF. It consists of 7 linear layers of width 128 and one output layer of width 64. All layers are activated by ReLU functions. A single skip connection concatenates the input with the fourth linear layer. $\gamma(\cdot)$ represents the positional encoding function.
In practice, we let the model predict the real and imaginary parts of the radiated signal, $(I,Q)$, instead of amplitude and phase, $(A,\psi)$. This is because phase is modulo against $2\pi$, which is not differentiable at some points. $A$ and $\psi$ can be easily translated from $I$ and $Q$ using Euler’s formula:
$A=\sqrt{I^{2}+Q^{2}},\;\psi=\arctan(Q/I)$
We use the same architecture for both "coarse" and "fine" models. A weighted sum of "coarse" and "fine" model loss is used for supervision:
@@
We use Adam optimizer Kingma and Ba [2014] with an initial learning rate of $5\times 10^{-4}$, and a ReduceLROnPlateau learning rate scheduler with patience 10 and factor 0.9. The other parameters are left as default. We train our model on a single NVIDIA A100 GPU, and it typically takes $\sim$100k iterations to converge.
The sampling range along each ray is adjusted for each environment due to their different dimensions. The sampling range is set to 9 m, 15 m, and 24 m for the conference room, bedroom, and office, respectively. For the conference room and bedroom, we take 128 samples per ray for the coarse stratified sampling, and another 128 samples per ray in the hierarchical sampling to enhance the granularity. For the office environment, due to its complexity, we double the number of samples for both sampling stages.

### v2

@@ 4.3 Improving the Convergence of NeWRF
One key difference between wireless measurements and images is that wireless antennas capture signals from all directions. Thus, in wireless, finding the ray direction to trace is a non-trivial problem. One potential solution is to search exhaustively along a discrete set of azimuth and elevation angle combinations that span all directions. However, this approach has two fatal drawbacks. First, the actual ray directions lie in a continuous space. Approximating them with a discrete set of ray directions would require an extremely dense grid of angles, which significantly increases the memory consumption for training the model. Secondly, for each receiver location, there are only 10$\sim$20 ray directions contributing significantly to the channel measurement, whereas a search grid of 1 resolution involves $360\times 180=64,800$ ray angles. As a result, optimizing the model to find the actual ray directions from such a large ray space is very difficult, if not impossible. We find through experiments that the model fails to converge with such a large search space of ray directions. To solve this issue, we leverage the fact that the Direction-of-Arrivals (DoA) of wireless signals can be measured using an antenna array at the receiver location (Foutz et al., 2022). We use DoA measurements as prior knowledge and let the learning algorithm trace rays primarily toward directions identified by the DoA measurements. Through our experiments, we find that this DoA-guided ray-casting approach significantly reduces the training time and memory consumption, as well as enables the learning algorithm to converge quickly.
@@ 4.4 Optimizing NeWRF
As in NeRF, We simultaneously optimize two MLP models for each environment. One ”coarse” and one ”fine”. We use mini-batch gradient descent to train our models where at each iteration, we randomly pick a batch of receiver locations from the training set and obtain the ground truth channel and DoA measurements. We cast rays toward each DoA and use the stratified sampling strategy (Mildenhall et al., 2021) to take coarse-grained samples uniformly distributed on each ray. We use these samples to query the ”coarse” model and use the procedures described in Section 4.2 to synthesize coarse predictions of the channel for each receiver location. We use the predicted weights $w_{\mathbf{p}}$ of coarse samples to perform a round of hierarchical sampling which densely samples regions with high weight values. We query the ”fine” model with the union of coarse and fine samples and synthesize the fine channel predictions. The coarse and fine channel predictions are each compared with the ground truth to calculate loss and update the parameters of each model.
In practice, we let the model predict the real and imaginary parts of the radiated signal, $(I,Q)$, instead of amplitude and phase, $(A,\psi)$. This is because phase is modulo against $2\pi$, which is not differentiable. $A$ and $\psi$ can be easily translated from $I$ and $Q$ using Euler’s formula:
$A=\sqrt{I^{2}+Q^{2}},\;\psi=\arctan(Q/I)$
@@
$L(\mathbf{h},\hat{\mathbf{h}})=\frac{\sum|\hat{\mathbf{h}}-\mathbf{h}|^{2}}{\sum|\mathbf{h}|^{2}}$ (8)
@@
Finally, it worth mentioning that similar to other NeRF work, we noticed the existence of floaters, i.e. spurious particles that ”float” in space, in the trained model. We find that casting a small number (5$\sim$10) of extra rays toward random directions during the training process can effectively regulate the model to remove the floaters, thus resulting in a clean geometry of the scene. Further details about our training pipeline and model architecture can be found in Appendix A.
@@ 4.5 NeWRF’s Scene Representation
Figure 6: The nature of wireless scenes: While backtracing direct rays leads to the transmitter, backtracing reflected rays consistently intersect at the virtual images of the transmitter.
We train the model as described in the previous section and find it has good generalization performance. However, we find that the model learns a representation of the wireless radiation scene that was different from what we originally expected. Despite that, the scene still allows the model to predict wireless fields (channels) even in room-scale environments with merely a sparse set of channel measurements. In this section, we present our findings and insights into it.
@@
To better understand why this is the case, we illustrate a simple wireless scene with a single transmitter and two reflective surfaces in Figure 6. Three receivers are placed in random locations to measure the wireless channel. The signals emitted by the transmitter propagate via three major paths to reach each receiver: one direct line-of-sight path and two reflection paths  [^3] We omit the second order reflection here for simplicity; however it can also be characterized in the same manner.. Since the receivers have no notion about the environment, they simply trace the rays back in the direction from which they arrived. Because of the law of reflection, the rays traced by all the receivers intersect at the locations of the virtual images of the transmitter and the transmitter itself. As a result, these points are sampled more frequently than other points in the space, giving the model an inductive bias to put higher volume densities at these locations.
We note that this representation is one of many possible solutions to this optimization problem that fits the training set. However, the model seems to have such a bias that it is more likely to converge to this representation of the scene. Although this representation is not what we had initially expected, it does reveal the simple nature of radiation scenes and allow for accurate channel synthesis from sparse measurements.
@@ 4.6 Predicting Channels at Unvisited Locations
While DoA information is known for our training data, it is not for the test data. Because of this, a new ray-casting technique was required for inference time predictions. Simply casting rays across a grid of all directions is problematic. Depending on the resolution of angles used for sampling we could under or overestimate the contribution of a virtual transmitter on a channel. When the resolution is too low, virtual transmitters could be missed and when it is too high, virtual transmitters could be sampled multiple times and contribute too much to the channel. For this reason, we developed a novel ray-searching algorithm that leverages our discovery of the nature of wireless scenes to extract DoA information and thereby intelligently cast rays at new locations.
Our ray-searching algorithm consists of five steps:
@@
    Intersection Point Identification: For each receiver location, $\mathbf{m}$, in the training set, perform ray marching along each DoA to identify intersection points, forming a set of potential (virtual) transmitters, $\mathbb{O}$.
@@
    Refine (Virtual) Transmitters: Apply DBSCAN (Ester et al., 1996) to cluster the estimated positions in $\mathbb{O}$. Add the centroid of each cluster to a reduced set, $\mathbb{V}$.
@@
    (Virtual) Transmitter Assignment: For each receiver location, $\mathbf{m}$, in the training set, find the closest (virtual) transmitter $\mathbf{v}\in\mathbb{V}$ along each DoA and add them to a set $\mathbb{V}_{m}$ of pertinent virtual transmitters for $\mathbf{m}$.
@@
    (Virtual) Transmitter Count Estimation for Testing Locations Train a fully connected neural network $g:\mathbf{m}\rightarrow\bigm{|}\mathbb{V}_{\mathbf{m}}\bigm{|}$ to map receiver locations to the corresponding number of (virtual) transmitters. Then use this model to predict the size of $\mathbb{V}_{\mathbf{t}}$, the set of pertinent (virtual) transmitters for test location $\mathbf{t}$, for all test locations.
@@
    Voting-based (Virtual) Transmitter Selection: For each $\mathbf{t}$, select its neighboring receiver locations from the training set (we use 6 nearest neighbors in practice). Populate $\mathbb{V}_{t}$ with the (virtual) transmitters most frequently assigned to the selected train locations. Finally, DoA can be calculated using the position of $\mathbf{t}$ and the estimated positions of (virtual) transmitters in $\mathbb{V}_{t}$.
We note that although this ray-searching algorithm is able to identify the locations of (virtual) transmitters, it cannot replace the role of training a NeWRF model for the scene, as we still need the prediction of signal amplitude and phase for each virtual transmitter.
@@ 5 Evaluation
To better understand the performance of NeWRF in representing wireless scene versus its performance in predicting channels in unseen locations, we first evaluate how accurately NeWRF can predict channels, assuming DoA is known for each receiver location. We then evaluate the performance of NeWRF in channel prediction while the DoA is not known and we use our ray-searching algorithm (presented in 4.6) to discover DoA of unseen locations.
Datasets: We generate simulation datasets for evaluation using MATLAB. We use the Image Method (Yun & Iskander, 2015) for ray tracing. We create datasets for three 3D environments models (MATLAB, 2023a, b; Free3D, 2023) of different complexities. Figure 7 shows the geometry and dimensions of the three models. We put a transmitter in each environment and randomly spread 443, 975, and 1907 receivers in each environment, respectively, to collect channel measurements. We use 80% of the measurements for training the model and the remaining 20% for testing.
Figure 8: Channel Prediction Results: a) Prediction results for conference room environment. Each dot represents the channel at one test location. b) Predicted channel components (paths) for a single location. Each dot represents a single propagation path. The arrow exhibits the sum of all paths; c) The channel prediction accuracy in different environments compared with baseline methods. d), e), and f) channel prediction accuracy for different training set sample densities, environment materials, and signal frequencies.
@@ 5.1 Channel Prediction Results
Figure 8(a) provides a visualization of the predicted channel for the test set measurements. Since the wireless channels are represented as complex numbers, we plot the real and imaginary parts of the channels in the x and y-axis, respectively. Each dot in the plot represents the channel for a particular receiver location. As is shown, our predictions align very well with the ground truth. Moreover, although we supervise the model with merely the sum of all multipath components, NeWRF can accurately predict individual path components for each particular receiver location as shown in Figure 8(b). In this plot, each dot represents one propagation path of the wireless signal; the arrow indicates the sum of all the components. Note that this capability of predicting individual path components is very important for optimizing the configurations of base stations and relays to enhance the network coverage in practice (Karanam & Mostofi, 2023).
@@
3) NeRF2: We use the open-sourced code of NeRF2 (Zhao et al., 2023) and train it on our datasets. Since their code assumes a dynamic transmitter scenario, we generate the counterpart simulation datasets by keeping the locations of each device the same as our datasets while swapping the roles of transmitters and receivers.
We use Signal-to-Noise Ratio (SNR), as our evaluation metric, which is defined as the SNR = $-10\log_{10}(\text{NMSE})$, where NMSE is Normalised Mean Square Error and defined in Eq. (8). Higher SNR indicates a higher prediction accuracy. For example, a 20 dB SNR is equivalent to a 1% relative squared error in the prediction. Figure 8(c) shows the results for three environments. NeWRF outperforms all of the baseline methods in all environments. Note, NeRF2 does not perform well on our datasets since the measurement density (0.4 $measurements/ft^{3}$) of our datasets is much lower than what they require (178 $measurement/ft^{3}$). Finally, our results show that although the performance of NeWRF degrades as the environment becomes larger and more complex (such as a large office space with lots of furniture), it still achieves more than 5 dB SNR which is sufficient for many wireless applications.
@@ 5.2 Measurement Density of Training Set
We further investigate the effect of training set sample density on the performance of NeWRF. To demonstrate this, we use the conference room dataset. In this experiment, we use 25%, 50%, 75%, and 100% of the training set measurements to train our model, which corresponds to 0.1, 0.2, 0.3, and 0.4 $measurements/ft^{3}$, respectively. We evaluate the trained model on the same test set for a fair comparison. As is shown in Figure 8(d), NeWRF can effectively learn the wireless radiation scene, and predict the channel accurately with sample density as low as 0.2 $measurements/ft^{3}$, which is around 800$\times$ lower than that of NeRF2(Zhao et al., 2023). This would significantly reduce the cost of wireless site surveys and improve the capability of dead spot detection.
@@ 5.3 Impact of Materials Types and Signal Frequencies
The attenuation and phase rotation of wireless signals during reflection are largely determined by the material that constitutes the reflector. In particular, the material’s electrical properties, such as permittivity and conductivity, play significant roles. As such, the wireless properties of an environment can change greatly as the materials it is made of change. Furthermore, even the same materials have different behaviors when they interact with signals of different frequencies. NeWRF is able to characterize both of these effects, as shown in Figure 8(e) and 8(f), where we train a separate model for each material type and signal frequency. NeWRF maintains high prediction accuracy in all these scenarios.
Figure 9: Ray-searching algorithm performance: Performance of NeWRF in channel prediction when DoA is known using ground truth or discovered using our Ray-searching algorithm. GT stands for ground truth, and RS stands for the Ray-Searching algorithm presented in  4.6.
5.4 Ray-searching Algorithm Performance
So far we evaluated the performance of NeWRF while DoA is known. Here, we evaluate the performance of NeWRF when we use our ray-searching algorithm to discover the DoA for each location. Figure 9 shows the results. Our results show that due to the error in DoA estimation, the performance of NeWRF in channel estimation degrades by a small margin (3$\sim$5 dB) when we use the ray-searching algorithm. This error is mainly due to a slight discrepancy of DoA between the test location and its nearest neighbors. Although this SNR loss is negligible, one can improve it by using slightly higher measurement density.
6 Limitations and Future work
In this work, we bring NeRF to radio frequency and demonstrate the promise of reconstructing wireless radiation scenes to facilitate site surveys. However, there remains a vast range of directions to investigate to unleash the power of NeRF for wireless applications. Here we highlight a few of them to help stimulate research in these directions:
a) Relaxing phase requirement: accurate phase measurement is challenging in real communication systems due to the synchronization error between the transmitter and receiver. The presented NeWRF algorithm, which relies on narrowband channels for training, does not perform well if the phase is inaccurate or not provided. One potential solution might be using wideband channel measurements for training since wideband channels translate phase into the variation of amplitude across different frequencies, which can be measured accurately.
b) Relaxing DoA requirement: accurate DoA measurement is required for ray-casting in the current NeWRF algorithm, the error tolerance is within one degree based on our experiments. In practice, achieving such a measurement resolution would require a very large antenna array. Further studies are needed to relax this constraint.
c) Extending to larger environments: although NeWRF outperforms all the baselines for the three environments studied, its performance reduction with the increasing environment complexity is observed. Scaling to more complex environments requires further research. One potential solution would be fusing visual and wireless radiation fields, where visual images provide more information about the environment to facilitate the learning of wireless fields.
@@ 7 Conclusion
Acknowledgements
We thank the UCLA ICON group and the reviewers for their insightful comments. We also thank Yadi Cao for his enlightening discussion. This work is partially supported by UCLA, and NSF awards 2238245 and 2146492.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Implementation Details
We implement NeWRF in Pytorch. Figure 11 shows the training pipeline. We use a batch size of 32 for each iteration of training. As in NeRF, we use positional encoding to map the input coordinates to higher dimensional space to facilitate the model to learn high-frequency variation of the wireless scene (Mildenhall et al., 2021). We use 10 frequencies to encode the spatial coordinates $(x,y,z)$, and 4 frequencies for the direction coordinates.
Figure 11 shows the MLP model architecture of NeWRF. It consists of 7 linear layers of width 128 and one output layer of width 64. All layers as well as the $\sigma$ output are activated by ReLU functions. The I, Q outputs are activated with an $tanh$ function to keep the output value within range [-1, 1]. A single skip connection concatenates the input with the fourth linear layer. $\gamma(\cdot)$ represents the positional encoding function.
We use the same architecture for both ”coarse” and ”fine” models. A weighted sum of ”coarse” and ”fine” model loss is used for supervision:
@@
We use Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of $5\times 10^{-4}$, and a ReduceLROnPlateau learning rate scheduler with patience 10 and factor 0.9. The other parameters are left as default. We train our model on a single NVIDIA A100 GPU, and it typically takes $\sim$100k iterations to converge.
The sampling ranges and resolutions along each ray are adjusted for each environment due to their different dimensions. The sampling range is set to 9 m, 15 m, and 24 m for the conference room, bedroom, and office, respectively. For the conference room and bedroom, we take 128 samples per ray for the coarse stratified sampling, and another 128 samples per ray in the hierarchical sampling to enhance the granularity. For the office environment, due to its complexity, we double the number of samples for both sampling stages.
Figure 12: NeWRF’s performance on adding noise to the channel
Figure 13: Subcarrier Channel Prediction